I really dont even know where to start. Translate the following into a quadratic equation, and solve it:
The length of a rectangular garden is three times its width; if the area of the garden is 75 square meters, what are its dimensions?

Answer by stanbon(75887)

The length of a rectangular garden is three times its width; if the area of the garden is 75 square meters, what are its dimensions?
Let the width be "x"
Then the length is "3x"
Area = length times width
75 (3x)x
25=x^2
x^2-25=0
(x-5)(x+5)=0
x=5 or x=-5
Only the answer x=5 is meaningful for your problem.
Cheers,
Stan H.